if given two functions f(x)=2x^2-5 and g(x)=3x-4 , find f(g(1))
Answer by jim_thompson5910(35256)

g(x)=3x-4

g(1)=3(1)-4

g(1)=3-4

g(1)=-1

-------------------------------------------------------

f(x)=2x^2-5

f(g(1))=2(g(1))^2-5

f(g(1))=2(-1)^2-5

f(g(1))=2(1)-5

f(g(1))=2-5

f(g(1))=-3